The sum of two numbers is 29. If you double the smaller number and the. Decrease it by the 10, the result is the larger number. Find the numbers
Answer by mananth(16946)

The sum of two numbers is 29. If you double the smaller number and the. Decrease it by the 10, the result is the larger number. Find the numbers
let the smaller number be x
the larger number be y
x+y = 29..........................1
2x-10=y
2x-y=10...........................2
add (1)& (2)
3x =39
/3
x=13
the other number = 39-13=16
the numbers are 13 & 16
...
m.ananth@hotmail.ca